Sol Summary – November 14th

Sol: 02
Summary Title: “Day Two, Seems Like Home”
Author’s name: Roger Gilbertson
Mission Status: Nominal
Sol Activity Summary: The morning was spent preparing and carrying out various individual projects. Cameras and equipment were setup for the afternoon EVA focusing on LIDAR and NeRF scanning. Science lab preparations continued for start of bacterial growth as soon as the pH meter arrives. Soil samples collected on EVA 2 were analyzed by crushing, sifting, then applying HCl and NaOH to small samples. Neither sample appeared to contain any appreciable amount of carbonates. GreenHab operations, planting and watering, continued with encouraging plant growth noticed.
A daily timeline was discussed and written up to streamline evening meal preparation and clean-up, to allow for more efficient use of the daily communications window, and earlier quiet and bed times. A step-by-step list was drafted for preparing, and returning from EVA. This includes using the three stations in the main downstairs room to setup each suit system in advance.
The afternoon EVA 3 departed about 30 minutes late and unfortunately the crew “assumed" a 30 minute extension without specifically requesting one from Mission Support. The EVA drove south on Cow Dung Road, noting a rapid drop in both vehicle’s state of charge. Communications became intermittent due to increasing distance from the repeater. One EVA crew member’s radio seemed to stop working, so they resorted to remaining close together in order to use gestures and shouting. CapCom was not notified of the radio problem until after the end of EVA, due to distance. The crew returned to the habitat at precisely 4:30 and entered the airlock at 4:35. Photos, documentary video, and 3D scans were successfully captured during the EVA. The crew will review the EVA requirements and communication protocols in the evening.
Look Ahead Plan: Tomorrow afternoon we plan for an EVA to collect soil samples in support of science started by a previous MDRS mission.
Anomalies in work: none
Weather: nominal
Crew Physical Status: nominal
EVA: EVA 3 completed today, request for EVA 4 submitted for tomorrow
Reports to be filed: Sol Summary, Journalist report, Photos, Operations report, Green Hab report, EVA 3 Summary, EVA 4 Request.
Support Requested: longest paint pole available to be attached to a window cleaning squeegee

Copyright © The Mars Society. All rights reserved. | Main Site